Sales volume last weekend for a sample of ten car dealerships were as follows:

Data Point 1 2 3 4 5 6 7 8 9 10
# Sales 23 13 37 46 20 15 26 45 33 38

What is the standard deviation of the sample data?

Please round your answer to the nearest tenth.

Note that the correct answer will be evaluated based on the full-precision result you would obtain using Excel.
Answer #1

А B с 1 2 х X^2 3 23 529 4 13 169 5 37 1369 6 2116 7 46 20 15 400 8 225 9 26 676 10 45 11 33 2025 1089 1444 38 296 10042 12 1

X dx = x-A = x - 30 dx2 23 -7 49 13 -17 289 37 7 49 46 16 256 20 -10 100 15 -15 225 26 -4 16 45 15 225 3 9 33 38 8 64 Σx = 29